Chapter 1 Systems of linear Equations
-Solve each of the linear systems of equations graphically. use a mini table of values to draw the graph of the lines.
x+y=2
3x=2y=9
so how would you graph both of these on one graph? and would you use the equation
Y= mx=b?

Answer by tutorcecilia(2152)

x+y=2 [Rewrite as y=mx+b by solving for y]
y=-x+2
Plug-in different values for the x's and solve for the y's.
Graph the results.
.
3x=2y=9 ?????[Rewrite as y=mx+b by solving for y]
If the equation is suppose to be:
3x-2y=9, than y = 3/2x-9/2
.
If the equation is suppose to be:
3x+2y =9, than y = -3/2x + 9/2
.
Plug-in different values for the x's and solve for the y's.
Graph the results.